OpenStudy (anonymous):

Evaluate each logarithmic expression: 1. (log2)16 + (log2)64 2. ln e^2

OpenStudy (anonymous):

\[1. \log_{2}16 + \log_{2}64 \]

myininaya (myininaya):

lne=1 lne^2=2lne=2(1)=2

OpenStudy (anonymous):

ln16/ln2=4 ln64/ln2=6 10 is the answer for the first one :) looks up for to turn log(basex)y into lny/lnx